To calculate the equilibrium concentration, you can use the ICE table approach. ICE stands for initial (1). change (C) and
Answer #1

Q1. (a.) From the calibration plot, the equation of straight line is : y = 4312x + 0.0075

where y-axis represents absorbance and x-axis represents equilibrium concentration of [FeSCN2+]

Given : absorbance = y = 0.32

From the equation of straight line,

x = (y - 0.0075) / 4312

x = (0.32 - 0.0075) / 4312

x = 7.25 x 10-5

equilibrium concentration of [FeSCN2+] = 7.25 x 10-5 M

(b.) initial concentration Fe3+ = 0.0010 M

initial concentration SCN- = 0.00060 M

ICE table [Fe3+] [SCN-] \rightleftharpoons [FeSCN2+]
I (initial) 0.0010 M 0.00060 M 0
C (change) -x -x +x
E (equilibrium) 0.0010 M - x 0.00060 M - x +x

Equilibrium concentration of [FeSCN2+] = 7.25 x 10-5 M (calculated) = x (from table)

Hence, x = 7.25 x 10-5 M

Equilibrium concentration Fe3+ = 0.0010 M - x

Equilibrium concentration Fe3+ = 0.0010 M - 7.25 x 10-5 M

Equilibrium concentration Fe3+ = 9.275 x 10-4 M

Equilibrium concentration SCN- = 0.00060 M - x

Equilibrium concentration SCN- = 0.00060 M - 7.25 x 10-5 M

Equilibrium concentration SCN- = 5.275 x 10-4 M

(c.) Equilibrium constant, K = [FeSCN2+]eq / [Fe3+]eq[SCN-]eq

K = (7.25 x 10-5 M) / [(9.275 x 10-4 M) * (5.275 x 10-4 M)]

K = 148.2